# FlowRL: Matching Reward Distributions for LLM Reasoning

### v1

@@ 3.2 FlowRL
4 Related Work
4.1 Reinforcement Learning for Reasoning
Reinforcement learning has emerged as a powerful approach for large language models post-training on reasoning tasks [Sutton et al., 1999b, Schulman et al., 2017, Lightman et al., 2023b, Shao et al., 2024, Guo et al., 2025]. Most approaches employ reward-maximizing RL to optimize expected cumulative returns. Entropy regularization [Haarnoja et al., 2018, Ahmed et al., 2019, Cheng et al., 2025] is a classical technique for mitigating mode collapse by promoting diversity in the policy’s output distribution, and has also been shown to enhance reasoning capabilities in various settings [Eysenbach and Levine, 2021, Chao et al., 2024]. However, for long CoT reasoning, the extended trajectory length (e.g., 8k–16k tokens) makes it difficult for the regularization signal to effectively influence reward-maximizing learning. Recent work [Cheng et al., 2025, Wang et al., 2025, Cui et al., 2025, Dong et al., 2025] has discovered that training with more diverse or high-entropy training data can further enhance training effectiveness. Compared to traditional entropy regularization, the above methods explicitly increase the proportion of low-probability (i.e., high-entropy) tokens in the training data. In our work, we address the mode-collapse problem by fundamentally shifting from reward maximization to reward distribution matching in our RL formulation.
4.2 GFlowNets
GFlowNets [Bengio et al., 2023a] represent a class of diversity-driven algorithms designed to balance probability flows across states. They have rich connections to probabilistic modeling methods [Zhang et al., 2022b, Zimmermann et al., 2022, Malkin et al., 2023, Zhang et al., 2022a, Ma et al., , Zhang et al., 2024a], and control methods [Pan et al., 2023c, d, b, Zhang et al., 2024b, Tiapkin et al., 2024]. This advantage has enabled GFlowNets to achieve successful applications in multiple downstream tasks, such as molecular drug discovery [Jain et al., 2022, 2023b, Liu et al., 2022, Jain et al., 2023a, Shen et al., 2023, Pan et al., 2023a, Kim et al., 2023, 2024], phylogenetic inference [Zhou et al., 2024], and combinatorial optimization [Zhang et al., 2023a, b]. For generative AI, GFlowNets provide a powerful approach to align pretrained models in scenarios such as image generation [Zhang et al., 2025a, Yun et al., 2025] and language model fine-tuning [Hu et al., 2024, Yu et al., 2025a, Lee et al., 2024]. Another line of work primarily focuses on the theoretical aspects of GFlowNets. Recent theoretical studies have interpreted GFlowNets as solving a maximum entropy reinforcement learning problem within a modified Markov Decision Process (MDP) [Tiapkin et al., 2024, Deleu et al., 2024, Mohammadpour et al., 2024]. These theoretical contributions have inspired us to enhance reinforcement learning from a more foundational standpoint using GFlowNets principles. A comprehensive overview of GFlowNets theory can be found in Appendix C.
4.3 Flow-Matching Policies
Flow matching simplifies diffusion-based approaches by learning vector fields that transport samples from prior to target distributions [Lipman et al., 2023]. Recent work has explored flow matching for policy optimization. McAllister et al. [2025] reformulates policy optimization using advantage-weighted ratios from conditional flow matching loss, enabling flow-based policy training without expensive likelihood computations. Pfrommer et al. [2025] explored reward-weighted flow matching for improving policies beyond demonstration performance. Park et al. [2025] uses a separate one-step policy to avoid unstable backpropagation through time when training flow policies with RL. Zhang et al. [2025a] proposed a combined loss function integrating PPO and GFlowNets to optimize diffusion model alignment. However, these approaches focus on continuous control, image generation, or vision-action models, rather than addressing mode-collapse limitations in reward-maximizing RL. Inspired by flow matching principles, our work improves upon RL training to enhance training stability while promoting diverse solution exploration.
@@ 5 Experiment Settings
There are two learnable modules in Eq. 6: the policy model $\pi_{\theta}$ and the partition function $Z_{\phi}$. For the policy model $\pi_{\theta}$, we use Qwen-2.5-7B/32B [Team, 2024] for math tasks and DeepSeek-R1-Distill-Qwen-7B [DeepSeek-AI, 2025] for code tasks, respectively. For partition function $Z_{\phi}$, following Lee et al. [2024], we use a randomly initialized 3-layer MLP with hidden dimensions matching those of the base model. The reference model $\pi_{\text{ref}}$ is the corresponding fixed pretrained model. All training scripts are based on the veRL [Sheng et al., 2024]. For the reward function, following Lee et al. [2024], we set the hyperparameter $\beta=15$.
@@
AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Qwen2.5-32B-Base, Max Response Len=8K
Backbone 4.6 2.1 28.6 52.5 27.0 21.4 22.7
R++ $14.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+10.2}}$ $9.2_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+7.1}}$ $52.7_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+24.1}}$ $44.4_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-8.1}}$ $17.4_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9.6}}$ $24.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.1}}$ 27.1
PPO $26.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+22.3}}$ $20.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+18.3}}$ $76.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+47.8}}$ $69.2_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+16.7}}$ $28.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+1.8}}$ $37.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+16.5}}$ 43.3
GRPO $23.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+18.5}}$ $14.6_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+12.5}}$ $76.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+48.3}}$ $61.6_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.1}}$ $19.0_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-8.0}}$ $34.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+13.5}}$ 38.3
FlowRL $24.0_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+19.4}}$ $21.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+19.8}}$ $73.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+45.2}}$ $80.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+28.3}}$ $38.2_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+11.2}}$ $51.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+30.4}}$ 48.4
Qwen2.5-7B-Base, Max Response Len=8K
Backbone 4.4 2.1 30.8 54.5 22.4 24.0 23.0
R++ $11.0_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.6}}$ $5.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.3}}$ $66.7_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+35.9}}$ $54.3_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.2}}$ $24.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.0}}$ $27.3_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.3}}$ 31.5
PPO $9.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.0}}$ $7.3_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.2}}$ $63.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+32.6}}$ $58.0_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.5}}$ $26.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+4.1}}$ $27.3_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.3}}$ 32.0
GRPO $13.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.1}}$ $9.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+7.7}}$ $64.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+33.7}}$ $57.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.6}}$ $23.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+0.7}}$ $26.9_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.9}}$ 32.5
FlowRL $15.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+11.0}}$ $10.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+8.7}}$ $54.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+23.7}}$ $67.0_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+12.5}}$ $31.4_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.0}}$ $34.6_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+10.6}}$ 35.6
@@
Table 1: Results on math benchmarks. We report Avg@16 accuracy with relative improvements shown as subscripts. Positive gains are shown in green and negative changes in red. FlowRL outperforms all baselines across both 7B and 32B model scales.
@@
DeepSeek-R1-Distill-Qwen-7B, Max Response Len=8K
Backbone 30.7 49.5 886.7 19.4 80.9
R++ $30.5_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.2}}$ $52.7_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.2}}$ $1208.0_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+321.3}}$ $56.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+37.4}}$ $76.6_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-4.3}}$
PPO $35.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+4.4}}$ $54.5_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.0}}$ $1403.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+516.4}}$ $73.7_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+54.3}}$ $82.3_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+1.4}}$
GRPO $32.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.1}}$ $52.3_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.8}}$ $1313.8_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+427.1}}$ $67.1_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+47.7}}$ $80.1_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.8}}$
FlowRL $\mathbf{37.4}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.7}}$ $\mathbf{56.3}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.8}}$ $\mathbf{1549.5}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+662.8}}$ $\mathbf{83.3}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+63.9}}$ $\mathbf{83.3}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.4}}$
Table 2: Results on code benchmarks. We report metrics with relative improvements shown as subscripts. Positive gains are shown in green and negative changes in red. FlowRL achieves the strongest performance across all three benchmarks, demonstrating its effectiveness in code reasoning tasks.

### v3

@@ 4 Experiment Settings
There are two learnable modules in Eq. 6: the policy model $\pi_{\theta}$ and the partition function $Z_{\phi}$. For the policy model $\pi_{\theta}$, we use Qwen-2.5-7B/32B [Team, 2024] for math tasks and DeepSeek-R1-Distill-Qwen-7B [DeepSeek-AI, 2025] for code tasks, respectively. The reference model $\pi_{\text{ref}}$ is the corresponding fixed pretrained model. For partition function $Z_{\phi}$, following Lee et al. [2024], we use a randomly initialized 3-layer MLP with hidden dimensions matching those of the base model. The input to $Z_{\phi}$ is the mean of the language model’s hidden states after encoding the input $\mathbf{x}$, and the output is a scalar value. We detail the implementation of $Z_{\phi}$ in § E. All training scripts are based on the veRL [Sheng et al., 2024]. For the reward function, following Lee et al. [2024], we set the hyperparameter $\beta=15$.
@@
Models AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Qwen2.5-32B-Base, Max Response Len = 8K tokens
Backbone 4.58 2.08 28.59 52.48 26.99 21.37 22.68
R++ $14.79_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+10.21}}$ $9.17_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+7.08}}$ $52.65_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+24.06}}$ $44.35_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-8.13}}$ $17.37_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-9.62}}$ $24.52_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.15}}$ 27.14
PPO $26.87_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+22.29}}$ $20.41_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+18.33}}$ $76.40_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+47.81}}$ $69.17_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+16.69}}$ $28.79_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+1.80}}$ $37.90_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+16.53}}$ 43.25
GRPO $23.12_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+18.54}}$ $14.58_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+12.50}}$ $76.87_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+48.28}}$ $61.60_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.12}}$ $18.95_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-8.04}}$ $34.94_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+13.57}}$ 38.34
FlowRL $23.95_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+19.37}}$ $21.87_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+19.79}}$ $73.75_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+45.16}}$ $80.75_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+28.27}}$ $38.21_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+11.22}}$ $51.83_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+30.46}}$ 48.39
Qwen2.5-7B-Base, Max Response Len = 8K tokens
Backbone 4.38 2.08 30.78 54.47 22.38 24.03 23.02
R++ $11.04_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.66}}$ $5.41_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.33}}$ $66.71_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+35.93}}$ $54.25_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.22}}$ $24.37_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+1.99}}$ $27.33_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.30}}$ 31.52
PPO $9.38_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.00}}$ $7.29_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.21}}$ $63.43_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+32.65}}$ $57.98_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.51}}$ $26.53_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+4.15}}$ $27.25_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.22}}$ 31.98
GRPO $13.54_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.16}}$ $9.79_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+7.71}}$ $64.53_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+33.75}}$ $57.05_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.58}}$ $23.06_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+0.68}}$ $26.88_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.85}}$ 32.48
FlowRL $15.41_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+11.03}}$ $10.83_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+8.75}}$ $54.53_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+23.75}}$ $66.96_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+12.49}}$ $31.41_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.03}}$ $34.61_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+10.58}}$ 35.63
@@
Table 1: Results on math reasoning benchmarks. We report Avg@16 accuracy with relative improvements shown as subscripts. Positive gains are shown in green and negative changes in red. FlowRL outperforms all baselines across both 7B and 32B model scales.
@@
DeepSeek-R1-Distill-Qwen-7B, Max Response Len = 8K tokens
Backbone 30.68 49.46 886.68 19.4% 80.90
R++ $30.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.22}}$ $52.68_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+3.22}}$ $1208.03_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+321.35}}$ $56.8\%_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+37.4\%}}$ $76.61_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-4.29}}$
PPO $35.10_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+4.42}}$ $54.48_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.02}}$ $1403.07_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+516.39}}$ $73.7\%_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+54.3\%}}$ $82.32_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+1.42}}$
GRPO $32.75_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.07}}$ $52.32_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.86}}$ $1313.82_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+427.14}}$ $67.1\%_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+47.7\%}}$ $80.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.77}}$
FlowRL $\mathbf{37.43}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.75}}$ $\mathbf{56.27}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+6.81}}$ $\mathbf{1549.47}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+662.79}}$ $\mathbf{83.3\%}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+63.9\%}}$ $\mathbf{83.28}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.38}}$
Table 2: Results on code benchmarks. We report metrics with relative improvements shown as subscripts. Positive gains are shown in green and negative changes in red. FlowRL achieves the strongest performance across all three benchmarks.